OpenStudy (anonymous):

Please help!! Factor the expression. 16m^2 – 24mn + 9n^2 A. (4m – 3n)(4m + 3n) B. (16m – 3n)(m + 3n) C. (4m – 3n)^2 D. (4m + 3n)^2

OpenStudy (anonymous):

its C

OpenStudy (anonymous):

thank u!

OpenStudy (anonymous):

(4m-3n)^2=16m^2+9n^2-24mn

OpenStudy (anonymous):

\[(a−b)^2=a^2−2ab+b^2\]